A block is released from rest at the top of an inclined 6.20 m long. The angle of the incline with respect to the horizontal direction is and the coefficient of kinetic friction between the block and the surfaces (incline and horizontal) is . The block slides along the incline with constant velocity and continues moving along the horizontal surface until it comes to rest. Using the work-energy theorem, Determine:

A block is released from rest at the top of an inc

a) The speed reached by the block at the bottom of the incline.   

b) The displacement of the block along the horizontal surface
Answer #1

if block was released from rest from the top position then it will acquire velocity as it comes down which can be calculated using work-energy theorem -

mgh = 0.5 mv2

v2 = 2gh

v2 = 2g*L cos theta

v2 = 2*9.8*6.20 cos theta

v2 = 121.52 cos theta

b> work done against friction willl be equal to kinetic energy of the block at the bottom-

u*mg*s = 0.5*m*121.52 cos theta

s = 6.2 cos theta/u

where u = coefficient of friction